Mix a solution that is 50% alcohol with a soution that is 20%to make 450 mL of a aolution that is 40% alcohol. How much of each solutions should you use ?
Answer by checkley71(8403)

.5X+.2(450-X)=.4*450
.5X+90-.2X=180
.5X-.2X=180-90
.3X=90
X=90/.3
X=300 ML OF 50% ALCOHOL.
450-300=150 ML OF 20% ALCOHOL
PROOF
.5*300+.2*150=180
150+30=180
180=180